OpenStudy (anonymous):

Given a triangle with a = 19, A = 43°, and B = 26°, what is the length of c? Round to the nearest tenth. I know that the final angle will be 21. But I don't know how I can find the side with just one side and 3 angles.

OpenStudy (jtvatsim):

the law of sines is your best bet.

OpenStudy (anonymous):

I'd need 2 sides to use the law of sines though :l

hero (hero):

Well, you have two angles and three add to 180, so find the missing angle first. Then use law of sines.

hero (hero):

43 + 26 + C = 180

OpenStudy (anonymous):

opps okai I added up to 90 for some reason >_<

hero (hero):

Okay, so @you-me-at-six, the correct value for C is...?

hero (hero):

I'm talking about the angle

OpenStudy (anonymous):

111?

hero (hero):

Okay, good. Now use law of sines

OpenStudy (anonymous):

I'm not sure how to apply it ._. (I learn all my stuff off of Youtube)

OpenStudy (anonymous):

Nevermind I tihnk I got....

hero (hero):

Let me know what you get

OpenStudy (agent0smith):

If you show your working, we can check you're using the law of sines correctly. \[\frac{ \sin A }{a } = \frac{ \sin B }{b} = \frac{ \sin C }{c }\]  Or you could look at this... but i felt like making a drawing, isn't it pretty? http://www.mathwarehouse.com/trigonometry/law-of-sines/images/formula-picture-law-of-sines2.png

hero (hero):

I get the feeling she may not get it the first time .

hero (hero):

But we'll see

OpenStudy (agent0smith):

@you-me-at-six drawing a picture might help you, label the sides/angles as in that drawing.

OpenStudy (anonymous):

I've been doing like different versions of this problem, but what I got was roughly 26 for length c

hero (hero):

Brilliantly done

OpenStudy (agent0smith):

Correct :)

OpenStudy (anonymous):

Thank you.

OpenStudy (agent0smith):

You're welcome. Especially for the pretty drawing.

OpenStudy (anonymous):

haha yes